OpenStudy (anonymous):

The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Write the linear equation for this problem in y=mx+b form. Note: this equation has a negative slope (m).

OpenStudy (rogue):

Y = 56 - x

OpenStudy (rogue):

The b is the initial temperature, your slope is -1 since the temperature is dropping 1 degree per hour. y = -1x + 56 = 56 -x

OpenStudy (anonymous):

how is the slope -1?

OpenStudy (anonymous):

sorry. i got it. " temperature is dropping 1 degree per hour"